2x + y = 7
x + 3y = 1
use substitution to solve

Answer by algebrahouse.com(1659)


2x + y = 7
x + 3y = 1 ------> x = -3y + 1 {subtracted 3y from both sides}

2(-3y + 1) + y = 7 {substituted -3y + 1, in for x, into top equation}
-6y + 2 + y = 7 {used distributive property}
-5y + 2 = 7 {combined like terms}
5y = 5 {subtracted 2 from both sides}
y = 1 {divided both sides by 5}

x = -3y + 1
x = -3(1) + 1 {substituted 1, in for y, into x = -3y + 1}
x = -3 + 1 {multiplied -3 by 1}
x = -2 {combined like terms}

x = -2 and y = 1
www.algebrahouse.com